OpenStudy (anonymous):

find the zeros of the function f(x) = 4x3 - 12x2 - 40x

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

not really the choices I have -5 and 2 5 and -2 0, 5, and -2 0, -5, and 2

OpenStudy (anonymous):

@skullpatrol Do you have this?

OpenStudy (skullpatrol):

@Paynesdad You can take it pal :)

OpenStudy (anonymous):

Laren the best bet is to try to factor if you can... always start by looking for common factors.

OpenStudy (skullpatrol):

$$\Huge 4x^3 - 12x^2 - 40x$$

OpenStudy (anonymous):

Remember to find zeros you set the function = 0 and solve for x ... right @lauren9695

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So the equation is \[4x^3-12x^2-40x=0\] Do you see any common factors in the left hand side?

OpenStudy (anonymous):

@lauren9695

OpenStudy (anonymous):

yes theyre all by a common factor of 4

OpenStudy (anonymous):

Ok and something else too

OpenStudy (anonymous):

ummm I cant think of anything else

OpenStudy (anonymous):

Well don't just think about numbers consider variables too.

OpenStudy (anonymous):

ohhh they have a variable of x

OpenStudy (anonymous):

Right so go ahead an factor the left side and show me what you have.

OpenStudy (anonymous):

@lauren9695

OpenStudy (anonymous):

wouldn't you the quadratic formula

OpenStudy (anonymous):

You can't yet because it is not a quadratic equation.... just factor the 4x out first and show me what you have ... we can go from there.

OpenStudy (anonymous):

@lauren9695 Hello...I will be leaving soon so I need you to respond a little quicker if you want my help.

OpenStudy (anonymous):

sorry my computer froze on me and if you factor out the 4x you would 3x^2+10x

OpenStudy (anonymous):

No there would still be three terms...What is left when you factor 4x^2 out of the 4x^3 term?

OpenStudy (anonymous):

oh 4x

OpenStudy (anonymous):

No if you factor 4x out of 4x^3 you will not be left with 4x ...try again.

OpenStudy (anonymous):

wouldn't it be x^3

OpenStudy (anonymous):

Not if you factored out an x too!

OpenStudy (skullpatrol):

$$\Huge 4x^3 - 12x^2 - 40x$$

OpenStudy (anonymous):

\[4x^3-12x^2-40x=4x\left(??-3x-10\right)\]

OpenStudy (anonymous):

Notice when I factored I didn't just lose or get rid of the 4x ... it just came out infront of the parentheses... but what is the ?? term @lauren9695

OpenStudy (skullpatrol):